OpenStudy (anonymous):

table x -2, 2 6 10 Y -4, -6, -8,-10 y=1/2x+5 or y=2x or y=-1/2x-5 or y=-2x-8 which equation best describes the relationship between x and y

OpenStudy (anonymous):

Just try them. y = 1/2(-2)+5 = 4, so that one is wrong for the very first pair. y = 2(-2) = -4, OK. y = 2(2) = 4, is wrong so that is not it. and so on for the last two equations, try them until they fail, if they don't fail for all the pairs then that's the one that best describes.

OpenStudy (anonymous):

thanks